# Effectiveness of the syndrome extraction circuit with flag qubits on IBM quantum hardware

### v1

@@ Acknowledgements
The sampled data are categorized into non-detected, corrected, and non-corrected. Non-detected samples are those in which the defect syndrome is not detected due to the absence of error. Corrected samples correspond to the data in which defect syndromes are detected and corrected; if these are not corrected, they belong to the non-corrected samples. In Fig. 24, the ratio of the categorized samples is shown with stacked bar charts for all cases of $[[d,1,d]]_{f}$ for both the hardware- and sample-based decoders. Although the samples are non-defective, a logical error can occur, which can be sorted as non-corrected instead of non-defective. As the structure size increases, the proportion of non-defects decreases. It is worth pointing out that most data samples are categorized as corrected samples as we increase the structure size.

### v2

@@ Acknowledgements
The sampled data are categorized into non-detected, corrected, and non-corrected. Non-detected samples are those in which the defect syndrome is not detected due to the absence of error. Corrected samples correspond to the data in which defect syndromes are detected and corrected; if these are not corrected, they belong to the non-corrected samples. In Fig. 24, the ratio of the categorized samples is shown with stacked bar charts for all cases of $[d,1,d]_{f}$ for both the hardware- and sample-based decoders. Although the samples are non-defective, a logical error can occur, which can be sorted as non-corrected instead of non-defective. As the structure size increases, the proportion of non-defects decreases. It is worth pointing out that most data samples are categorized as corrected samples as we increase the structure size.